Use six unit multipliers to convert 30 cubic yards to cubic inches?
Answer by CubeyThePenguin(3113)

30 yd^3 = 30 * 3^3 ft^3 = 810 ft^3

810 ft^3 = 810 * 12^3 in^3 = 1399680 in^3